OpenStudy (anonymous):

Write one example of a linear function and one example of a non-linear function. (Use x and y as the variables)

OpenStudy (anonymous):

@Lovelarap @the_ocean_girl @Mehek14

OpenStudy (anonymous):

help how do I do this?

OpenStudy (anonymous):

@IrishBoy123 @jim_thompson5910

OpenStudy (anonymous):

@micahm @mathmate

OpenStudy (anonymous):

A linear function has the following form. y = f(x) = a + bx.

OpenStudy (anonymous):

non liner is not a line for example a parabola.

OpenStudy (anonymous):

lol yea ik and there are ways you can tell its non linear

OpenStudy (anonymous):

a liner is simply a line. ; ) figured it out

OpenStudy (anonymous):

haha yeah

OpenStudy (anonymous):

so i need an example like in numbers hmm... maybe i can fugure it out

OpenStudy (anonymous):

a parabola has a x^2, liner functions are simple x

OpenStudy (dayakar):

f(x,y) =ax+by+c [where a,b c are real numbers]

OpenStudy (dayakar):

non linear function u can write any power on variables other than 1

OpenStudy (anonymous):

thx guys i figured it out lemme post what im putting wait a sec plz

OpenStudy (anonymous):

An example of a linear function would be y = 2x+1 because on a graph it turns out to be a straight line. An example of a non - linear function would be 4x2 + 2y - 1 = 0 because on a graph it would not look like a line so since it isn't like a straight line, which basically describes a linear function, then it is a NON - linear function, it is NOT linear.

OpenStudy (dayakar):

correct

OpenStudy (anonymous):

ugh why did it send that again

OpenStudy (anonymous):

nevermind thanks again!!! :D